A circular metal plate with radius 0.22 m. Calculate the net torque about O given that F1 = 17.4 N, F2 = 7.2 N, F3 = 14 N, F4 = 5.4 N, θ1 = 58°, and θ2 = 38°.





1
Expert's answer
2021-03-22T08:36:41-0400

"F_2\cdot\sin38°\cdot0.22-F_1\cdot\sin58°\cdot0.22+F_4\cdot0+F_3\cdot0.22="


"=7.2\cdot\sin38°\cdot0.22-17.4\cdot\sin58°\cdot0.22+"


"+5.4\cdot0+14\cdot0.22=0.81\ (N\cdot m)" clockwise . Answer